OpenStudy (anonymous):

ill medal and its multiple choice . During a car's lifetime its value will slowly decrease and then, once it becomes a classic, will slowly increase. If the value, V(t) is a function of time, then during part of a particular car's life V(t) = 15.4t2 - 785t + 12000. According to this model, what will be the car's lowest approximate value? $1500 $2000 $2500 $3000

OpenStudy (misty1212):

\[V(t) = 15.4t^2 - 785t + 12000\] you want the vertex of this parabola

OpenStudy (misty1212):

the second coordinate of the vertex really the first coordinate is \(-\frac{b}{2a}\) and the second coordinate is what you get when you replace \(x\) by that number

OpenStudy (misty1212):

in your example \[a=15.4,b=-785\] find \[\frac{785}{2\times 15.4}\] first

OpenStudy (solomonzelman):

mmmm, I would first factor out of the coefficient of x^2, then be adding inside the parenthesis. Such. \(\large\color{black}{ y=c_1(x^2+c_2x)+c_3 }\) \(\large\color{black}{ y=c_1(x^2+c_2x\color{red}{+(c_2/2)^2-(c_2/2)^2})+c_3 }\) \(\large\color{black}{ y=c_1(x^2+c_2x\color{red}{+(c_2/2)^2})+c_3 \color{red}{-\color{black}{c_1}(c_2/2)^2} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ c_1,~~c_2,~~c_3 }\) are just 3 different numbers.

OpenStudy (anonymous):

thats confusing.

OpenStudy (misty1212):

wow that is an awful lot of work to compute the first coordinate of the vertex \(-\frac{b}{2a}\) i guess you could also take the derivative, set it equal to zero and solve probably get \(-\frac{b}{2a}\) also

OpenStudy (solomonzelman):

no, taking the derivative might be something that the user haven't learned yet. Also this actually isn't hard.

OpenStudy (solomonzelman):

maybe the way I put it with c1 and c2, but we can do an example, if you want.

OpenStudy (anonymous):

yes pleas.

OpenStudy (anonymous):

is it $2000 @SolomonZelman

OpenStudy (misty1212):

there may be a typo in your question the minimum value of that parabola is somewhere around \(-8803.65\)

OpenStudy (misty1212):

oops no you are right, thanks @freckles

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=15.4t^2+-+785t+%2B+12000 \[1996.35\]